# The CAVITY project. The spatially resolved stellar population properties of galaxies in voids

### v1

@@ 1 Introduction
The Calar Alto Void Integral-field Treasury surveY (CAVITY, Pérez et al. in preparation) is a legacy project of the Calar Alto Observatory. It aims to observe the first statistically complete IFU data set of galaxies in voids, making use of the PMAS/PPaK instrument located at the 3.5m telescope of Calar Alto (the same as in CALIFA). The main goals of this project consist in studying the mass assembly (baryonic and dark) and the gas and stellar properties of galaxies in voids to discover the effect of the low density environment on the formation and evolution of galaxies.
@@
This paper is organised as follows. For the sake of clarity, we summarise in Sections 2 and 3 the observations, data reduction, and sample selection, which are explained in detail in the CAVITY presentation paper (Pérez et al, in prep.). The methodology and the calculations to obtain the stellar population properties are then presented in Section 4. In Section 5 we present our main results, and we compare them with a control sample of galaxies in filaments and walls in Section 6. These results are discussed in Section 7. Finally, Section 8 summarises the work and the conclusions. The masses calculated for this work were normalised with a Salpeter initial mass function (IMF). Throughout this work we assume a flat $\Lambda$CDM (Lambda cold dark matter) cosmology with $\Omega_{M}=0.3$, $\Omega_{\Lambda}=0.7$, and $H_{0}=70$ km/s/Mpc.
@@ 4.2 Stellar population properties
We determine the current SFR by summing the stellar mass formed over the latest time interval and dividing it by its duration (Equation 3). We define a star-forming age, $t_{\text{SF}}$ (the age of the older stars taken into account in the sum), as an arbitrary value for this calculation.
@@ 5.1 Stellar mass
The distribution of stellar masses (M), obtained from the spectral fitting, for each morphological type is shown in Figure 4. The mass range between $1.0\times 10^{9}$ and $4.2\times 10^{11}\,$, having its median at $2.6\times 10^{10}\,$, and it can be seen that they are slightly correlated with the Hubble type as expected, given that we are not including dwarf galaxies: earlier types tend to be more massive than later types. Sbs have the largest dispersion, with values ranging from $10^{9}$ to $10^{11.3}\,$.
@@ 5.3 Stellar ages
Figure 8 shows the age radial profiles, stacked by morphology (upper panel) and by stellar mass and morphology (bottom panel). They show the expected decrease of the age from the centre towards larger distances, reflecting the inside-out formation scenario of galaxies (e.g. Pérez et al. 2013; Sánchez-Blázquez et al. 2014; González Delgado et al. 2014, 2017; García-Benito et al. 2017; Goddard et al. 2017). It is worth noting that this scenario is not universal for all kinds of galaxies, but is restricted to a certain types and stellar mass ranges. Very massive early type galaxies ($\text{M}\sim 10^{12}\,\text{M}_{\odot}$), usually those in the centre of clusters, are found to have flat age profiles (Oliva-Altamirano et al., 2015), while dwarf ellipticals ($\text{M}<10^{9}\,\text{M}_{\odot}$) tend to have positive age gradients (Bidaran et al., 2023). However, none of the galaxies analysed in this work lay on any of these groups.
@@ 6.4 Radial profiles
We compare the radial structure of the stellar population properties of galaxies in voids and in filaments and walls, to determine if the differences identified in the integrated properties are occurring everywhere in the galaxies, or instead, they are specifically associated with the inner regions (R ¡ 1 HLR) or outskirts (R ¿ 1 HLR) of the galaxies. While the light from the central part of S0 and spirals is probably more associated with the bulge, the light from regions located at distances larger than 1 HLR comes from the discs. Thus, finding out where the variations come from can give us clues about different formation scenarios and/or evolution of the bulges and discs of galaxies in low-density environments.
@@
Figure 17 shows the radial profiles of the properties we are analysing. In each panel, we represent the stacked profiles of each morphological type of the void sample and the filament and wall sample. Like before, we decided not to plot the Sds, as we only have one.
@@
In the case of the light-weighted age, on the panels of the second row, we found the opposite: there is no significant difference in the earlier types, but in the spirals, where we find a mismatch in the outer parts. We find the biggest difference from 0.5 HLR outwards, outside of what could be the bulge of the spiral types. The discs of spiral void galaxies seem to be less evolved, with more pronounced gradients than those from galaxies in filaments and walls. In the Sas, we see differences not only in the outer parts but at every galactocentric radius except for the central region.
@@ 7.2 The quenching of star formation in voids
Quenching is caused by a combination of these effects. It is known that the fraction of galaxies that have undergone this transformation is higher in environments of high density (George et al., 2011), specially near the centres of clusters (Donnari et al., 2021). Furthermore, the fraction of quenched galaxies in groups increases with galaxy stellar mass, but is always higher in groups than in the field (González Delgado et al., 2022). We want to test if a similar difference is also found for the environments of least number-density with respect to filaments and walls.
@@ 7.3 In situ star formation and gas accretion mode in void galaxies
Differences in the star formation properties may also be due to changes in the relation between baryonic and dark matter in different environments. Works based on simulations (Alfaro et al., 2020; Habouzit et al., 2020; Rosas-Guevara et al., 2022; Rodríguez-Medrano et al., 2023), find that for a fixed stellar mass, dark matter haloes are more massive in voids than in denser environments. This means that for a fixed stellar mass, the gravitational potential well in which galaxies evolve is higher in voids, which can affect their properties. It has been shown that the mass of the haloes affect the SFHs of the galaxies that inhabit them (Scholz-Díaz et al., 2023), and that higher halo mass-to-stellar mass ratios result in younger stellar ages (Scholz-Díaz et al., 2022), which is consistent with our results.
@@ 7.4 Mass assembly, ex situ accretion of stars, and mergers in voids
It is easier to trace the mass accretion histories through simulations than with observed data. They do so in Rodríguez-Medrano et al. (2023), using the Illustris TNG300 simulation (Nelson et al., 2019; Pillepich et al., 2018). They find that the mass accretion has happened more recently in voids than in other environments. Mergers (ex situ processes) are one of the main drivers of mass accretion in galaxies. It has been shown that although the number of mergers does not vary between different environments, they happen at later epochs in voids with respect to denser environments (Rosas-Guevara et al., 2022; Rodríguez-Medrano et al., 2023).
In this work, we compare the stellar mass surface density profiles of the galaxies in cosmic voids to those of our control sample, and find that globally, galaxies in voids are less dense than their counterparts. This happens specially for ellipticals, although we must remember that it is also this morphological type that has the biggest shift in stellar mass. One could attribute this shift in the stellar mass density to mergers being more common in filaments and walls (specifically, dry mergers, as the age of the morphological types that are more different in density are not in age). However, as it was already discussed, the number of mergers in both environments has been found to be equivalent.
@@ 8 Summary and conclusions
  • The SFR is slightly higher in galaxies than in voids, by $\Delta_{\text{void-fil}}$(log SFR) = $0.15\pm 0.72$, specially in early types. However, we do not find this difference to be statistically significant enough except for Sas. The amount of quenched ETG in filaments and walls is higher, indicating a slower transition from star-forming to quiescent in void galaxies.
@@ Appendix A Table of global properties
Table 5: Stellar population and structural properties of each CAVITY galaxy in our sample.

### v3

@@ 1 Introduction
The Calar Alto Void Integral-field Treasury surveY (CAVITY, Pérez et al. 2024) is a legacy project of the Calar Alto Observatory. It aims to observe the first statistically complete IFU data set of galaxies in voids, making use of the PMAS/PPaK instrument located at the 3.5m telescope of Calar Alto (the same as in CALIFA). The main goals of this project consist in studying the mass assembly (baryonic and dark) and the gas and stellar properties of galaxies in voids to discover the effect of the low density environment on the formation and evolution of galaxies.
@@
This paper is organised as follows. For the sake of clarity, we summarise in Sections 2 and 3 the observations, data reduction, and sample selection, which are explained in detail in the CAVITY presentation paper (Pérez et al., 2024). The methodology and the calculations to obtain the stellar population properties are then presented in Section 4. In Section 5 we present our main results, and we compare them with a control sample of galaxies in filaments and walls in Section 6. These results are discussed in Section 7. Finally, Section 8 summarises the work and the conclusions. The masses calculated for this work were normalised with a Salpeter initial mass function (IMF). Throughout this work we assume a flat $\Lambda$CDM (Lambda cold dark matter) cosmology with $\Omega_{M}=0.3$, $\Omega_{\Lambda}=0.7$, and $H_{0}=70$ km/s/Mpc.
@@ 4.2 Stellar population properties
We determine the current SFR by summing the stellar mass formed over the latest time interval and dividing it by its duration (see Equation 3). We define a star-forming age, $t_{\text{SF}}$ (the age of the older stars taken into account in the sum), as an arbitrary value for this calculation.
@@ 5.1 Stellar mass
The distribution of stellar masses (M), obtained from the spectral fitting, for each morphological type is shown in Figure 4. The mass ranges between $1.0\times 10^{9}$ and $4.2\times 10^{11}\,$, having its median at $2.6\times 10^{10}\,$, and it can be seen that it is slightly correlated with the Hubble type as expected, given that we are not including dwarf galaxies: earlier types tend to be more massive than later types. Sbs have the largest dispersion, with values ranging from $10^{9}$ to $10^{11.3}\,$.
@@ 5.3 Stellar ages
Figure 8 shows the age radial profiles, stacked by morphology (upper panel) and by stellar mass and morphology (bottom panel). They show the expected decrease of the age from the centre towards larger distances, reflecting the inside-out formation scenario of galaxies (e.g. Pérez et al. 2013; Sánchez-Blázquez et al. 2014; González Delgado et al. 2014, 2017; García-Benito et al. 2017; Goddard et al. 2017). It is worth noting that this scenario is not universal for all kinds of galaxies, but is restricted to certain types and stellar mass ranges. Very massive early type galaxies ($\text{M}\sim 10^{12}\,\text{M}_{\odot}$), usually those in the centre of clusters, are found to have flat age profiles (Oliva-Altamirano et al., 2015), while dwarf ellipticals ($\text{M}<10^{9}\,\text{M}_{\odot}$) tend to have positive age gradients (Bidaran et al., 2023). However, none of the galaxies analysed in this work lay on any of these groups.
@@ 6.4 Radial profiles
We compare the radial structure of the stellar population properties of galaxies in voids and in filaments and walls, to determine if the differences identified in the global properties are occurring everywhere in the galaxies, or instead, they are specifically associated with the inner regions (R ¡ 1 HLR) or outskirts (R ¿ 1 HLR) of the galaxies. While the light from the central part of S0s and spirals is probably more associated with the bulge, the light from regions located at distances larger than 1 HLR comes from the discs. Thus, finding out where the variations come from can give us clues about different formation scenarios and/or evolution of the bulges and discs of galaxies in low-density environments.
@@
Figure 17 shows the radial profiles of the properties we are analysing. In each panel, we represent the stacked profiles of each morphological type of the void sample and the filament and wall sample. Like before, we decided not to plot the Sd, as we only have one.
@@
In the case of the light-weighted age, on the panels of the second row, we find the opposite: there is no significant difference in the earlier types, but in the spirals, where we see a mismatch in the outer parts. We find the biggest difference from 0.5 HLR outwards, outside of what could be the bulge of the spiral types. The discs of spiral void galaxies seem to be less evolved, with more pronounced gradients than those from galaxies in filaments and walls. In the Sas, we see differences not only in the outer parts but at every galactocentric radius except for the central region.
@@ 7.2 The quenching of star formation in voids
Quenching is caused by a combination of these effects. It is known that the fraction of galaxies that have undergone this transformation is higher in environments of high density (George et al., 2011), especially near the centres of clusters (Donnari et al., 2021). Furthermore, the fraction of quenched galaxies in groups increases with galaxy stellar mass, but is always higher in groups than in the field (González Delgado et al., 2022). We want to test if a similar difference is also found for the environments of least number-density with respect to filaments and walls.
@@ 7.3 In situ star formation and gas accretion mode in void galaxies
Differences in the star formation properties may also be due to changes in the relation between baryonic and dark matter in different environments. Works based on simulations (Alfaro et al., 2020; Habouzit et al., 2020; Rosas-Guevara et al., 2022; Rodríguez-Medrano et al., 2024), find that for a fixed stellar mass, dark matter haloes are more massive in voids than in denser environments. This means that for a fixed stellar mass, the gravitational potential well in which galaxies evolve is higher in voids, which can affect their properties. It has been shown that the mass of the haloes affect the SFHs of the galaxies that inhabit them (Scholz-Díaz et al., 2023), and that higher halo mass-to-stellar mass ratios result in younger stellar ages (Scholz-Díaz et al., 2022), which is consistent with our results.
@@ 7.4 Mass assembly, ex situ accretion of stars, and mergers in voids
It is easier to trace the mass accretion histories through simulations than with observed data. They do so in Rodríguez-Medrano et al. (2024), using the Illustris TNG300 simulation (Nelson et al., 2019; Pillepich et al., 2018). They find that the mass accretion has happened more recently in voids than in other environments. Mergers (ex situ processes) are one of the main drivers of mass accretion in galaxies. It has been shown that although the number of mergers does not vary between different environments, they happen at later epochs in voids with respect to denser environments (Rosas-Guevara et al., 2022; Rodríguez-Medrano et al., 2024).
In this work, we compare the stellar mass surface density profiles of the galaxies in cosmic voids to those of our control sample, and find that globally, galaxies in voids are less dense than their counterparts. This happens especially for ellipticals, although we must remember that it is also this morphological type that has the biggest shift in stellar mass. One could attribute this shift in the stellar mass density to mergers being more common in filaments and walls (specifically, dry mergers, as the morphological types that are more different in density are not in age). However, as it was already discussed, the number of mergers in both environments has been found to be equivalent.
@@ 8 Summary and conclusions
  • The SFR is slightly higher in galaxies than in voids, by $\Delta_{\text{void-fil}}$(log SFR) = $0.15\pm 0.72$, especially in early types. However, we do not find this difference to be statistically significant enough except for Sas. The amount of quenched ETG in filaments and walls is higher, indicating a slower transition from star-forming to quiescent in void galaxies.
@@ Appendix A Table of global properties
Table 5: Global stellar population and structural properties of each CAVITY galaxy in our sample.